Circular No. 7378

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVA 2000P IN NGC 4965 F. Colas, Institut de Mécanique Céleste, Paris, and Observatoire du

Pic du Midi, reports the discovery by Robin Chassagne (Sainte Clotilde, Île de Réunion) of an apparent supernova (unfiltered mag 14.1) on a 180-s exposure taken with a 0.30-m telescope (+ ST-7E CCD camera) on Mar. 8.03 UT. Colas confirmed the presence of SN 2000P (R = 14.5) with the 1.05-m telescope at Pic du Midi on Mar. 9.12, noting that the new object does not appear on the Digital Sky Survey (limiting mag ∼ 21); he measured the position of SN 2000P as α = 13^{h}07^{m}10^{s}.53, δ = −28^{o}14′02′′.5 (equinox 2000.0), which is 16′′ east and 21′′ south of the center of NGC 4965. Chassagne measured position end figures 10^{s}.56, 02′′.6. H. Yamaoka, Kyushu University, also reports the following position end figures measured from the discovery image: 09^{s}.88, 13′59′′.3.

SUPERNOVA 2000O IN MCG +3-31-61 S. Jha, P. Challis, and R. Kirshner, Harvard-Smithsonian Center for

Astrophysics, report that a spectrum of SN 2000O (cf. IAUC 7375) was obtained by Jha on Mar. 8.4 UT with the F. L. Whipple Observatory 1.5m telescope (+ FAST spectrograph), showing it to be a type-Ia supernova after maximum light. The spectral-feature age for the supernova is 13 ± 4 days past maximum (Riess et al. 1997, A.J. 114, 722). The Si II (rest 635.5 nm) expansion velocity is 10 200 km/s, adopting the NED recession velocity of 7042 km/s for the host galaxy.

E. Cappellaro, S. Benetti, M. Turatto, A. Pastorello, S. Ortolani, and E. M. Corsini, Padua, report that a spectrum (range 384–684 nm, resolution 1.0 nm) of SN 2000O, obtained on Mar. 9.3 UT with the Danish 1.54-m telescope (+ DFOSC) at the European Southern Observatory, is very similar to that of the type-Ia supernova 1994D (cf. Patat et al. 1996, MNRAS 278, 111) two weeks after maximum light. The expansion velocity of ∼ 10 900 km/s, deduced from the minimum of Si II 635.5-nm is consistent with the estimated epoch (the recession velocity of 6900 km/s has been measured from the Na I D insterstellar absorptions, with equivalent width ∼ 0.14 nm).

CORRIGENDA On IAUC 7370, ‘Comet P/1999 XB_{69} (LINEAR)’, line 5, for q =

2000 Feb. read T = 2000 Feb. Also on IAUC 7370, ‘Comet P/1999 XN_{120} (Catalina)’, line 4, for q = 2000 May read T = 2000 May

2000 March 10  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2000 CBAT  Daniel W. E. Green